# Forward modeling fluctuations in the DESI LRGs target sample using image simulations

### v1

@@ 7.2 Galactic Extinction
Galactic extinction is caused by dust grains in the Milky Way. The Milky Way dust scatters light from distant galaxies, making them appear dimmer to us. Blue light is scattered more strongly than red light. The thickness of dust varies across the sky, introducing an inhomogeneous reddening effect. Various maps have been developed to correct for this effect. As detailed in [68], these maps can be divided into three categories: Thermal emission-based maps, e.g., [65], stellar reddening-based maps, e.g., [69], and HI-based maps, e.g., [70]. The Galactic extinction is typically mapped as $E(B-V)$, in magnitudes, and a correction to apply to any measured magnitude is a constant coefficient for the particular wavelength band, multiplied by the $E(B-V)$ at the given celestial coordinate. The SFD ([65]) map is the most commonly used map for the application of extinction corrections in current galaxy surveys, including the DESI LRGs targets. It is thermal-based, and the data are directly collected from dust emission. This gives it high accuracy. However, it suffers from extragalactic imprints. When stacking the SFD $E(B-V)$ value around reference objects, [68] see that the average $E(B-V)$ value correlates with the redshift of these objects, indicating leaks of Cosmic Infrared Background (CIB) signal. The stellar-reddening-based maps measure dust emission on a different wavelength range, in which CIB contamination is negligible. These maps measure $E(B-V)$ by comparing stellar spectra to templates while fitting for the $E(B-V)$ value at the same location. This method is less accurate than the thermal emission-based map, with a significant limitation being the number of usable stars at any given location, but is free from CIB leaks. HI-based maps trace the hydrogen column density, which is correlated with $E(B-V)$. HI-based maps are not strictly $E(B-V)$ maps. Due to the scope of this work, we did not use HI-based maps in our test.
Observations of distant galaxies are affected by $E(B-V)$. If we have a “truth map” for $E(B-V)$ (and per-band coefficients), it would not be a big problem, as the needed corrections to the photometry would be known. However, we do not have such a “truth map”. When we detect a correlation of our galaxies’ relative density with the $E(B-V)$ map, we are not certain how the error in our $E(B-V)$ map plays a role in such correlation. We will consider three potential factors that could contribute to the trends we observe between $E(B-V)$ and DESI LRGs:
@@
    The $E(B-V)$ values are correlated with the true background galaxy density due to some Large-Scale-Structure systematics in the $E(B-V)$ map, e.g., the CIB leaks. For LRGs, their own dust can absorb light and re-emit it in the infrared, contributing to the CIB. Thus, regions with a higher density of LRGs will have higher CIB and a positive trend should be expected between the LRG density and an $E(B-V)$ map with CIB contamination.
@@
Factors 1 and 2 are not recovered by Obiwan because Obiwan assumes no error in the $E(B-V)$ map, while factor 3 should be automatically included during Obiwan’s image simulation procedure. In DESI LRGs, we observe a strong correlation of the relative number density of LRGs with the $E(B-V)$ SFD map, as shown in Figures 12 and 13. Meanwhile, Obiwan LRGs exhibit a weaker correlation. Under our assumption, the weaker correlation in Obiwan LRGs is due to not simulating factors 1 and  2 in the Obiwan simulation.
To test the impact of the three factors on the LRG density vs. $E(B-V)$ trend, we redo the DESI LRGs target selection based on two alternative maps. One of the alternative maps is the stellar-reddening-based map produced by [69], which we label ‘Mudur+23’. This map derives $E(B-V)$ map based on data from Pan-STARRS1 [71] and 2MASS [72] along with Gaia parallaxes [73]. The other alternative map is the Corrected-SFD (CSFD) map produced by [74]. It cross-correlated the SFD map with the galaxy clustering signal and separated the SFD map into two maps: the CIB residual map and the CSFD map. The CSFD dust map has a much lower residual CIB contamination than SFD. However, the correction is CIB only. Non-CIB Large-Scale-Structure systematics are not included in the CSFD map. Compared to the stellar reddening-based map, the $E(B-V)$ values on the CSFD map share more similarity with the SFD map. We define LRGs selected with these maps real LRGs, since they are all real galaxies we observe. When using an alternative $E(B-V)$ map, real LRGs is not always identical to DESI LRGs. When we use the alternative $E(B-V)$ maps, the (extinction-corrected) magnitudes of the galaxies change. When we apply LRG target selection to these galaxies, we obtain a sample that is slightly different from DESI LRGs. Meanwhile, if we use SFD map in LRG target selection, real LRGs are identical to DESI LRGs.
@@
Figure 18: Density fluctuation of DESI LRGs based on different maps in the NGC (left) and SGC (right) region. The target selection process is also altered based on different $E(B-V)$ maps. The blue line is based on the SFD map [65], the green line (labeled ‘Mudur+23’) is based on the stellar reddening map from [69], the orange line is based on the Corrected-SFD (CSFD) map [74], the red line uses LRGs selected by the SFD map, but the binning is computed with the same stellar reddening map as the green line. For LRGs selected and plotted against SFD maps, we see a clear trend. In the stellar reddening map, NGC and SGC are slightly different. In NGC, we do not see any trend. In SGC, we see a trend that is weaker than the SFD map. The difference is also seen in Obiwan LRGs seen in Fig. 19. In the CSFD map, this trend is similar to the SFD map. In the lower $E(B-V)$ region, the trend is weaker. The fact that the red lines exhibit a larger trend than the green line suggests that some of the $E(B-V)$ systematics trend is introduced by target selection, which uses the $E(B-V)$ to determine the colors of galaxies.
Factor 1 can be seen by comparing the red curve with the green curve. The two samples are binned with the same stellar-reddening-based map. The red sample is DESI LRGs that uses SFD map to determine their magnitudes and perform target selection. The green curve uses a stellar-reddening-based map for a similar procedure. Thus, the difference between these two curves is a result of the change in target selection. We find that 2.0% of LRGs are selected with the stellar-reddening based map, and not selected by the SFD map. 2.6% of LRGs are selected with the SFD map, but not selected by the stellar-reddening based map. These two curves in Figure 18 show how LRG sample difference results in different systematics trends. The red curve has a larger trend than the green curve, which indicates that the CIB leaks affect the target selection of DESI LRGs.
Factor 2 can be seen by comparing the blue curve and the red curve. The two curves are computed with identical DESI LRGs sample. The only difference lies in the binning. The blue curve is binned with the SFD map, and has a stronger trend. The red curve is binned with the stellar-reddening-based map, and has a weaker trend. The reduction in trend in the red curve compared to the blue curve is a result of less CIB contamination in the stellar-reddening-based map. If we assume the Mudur+23 map is a “truth map”, then the excess trend in the blue curve emerges by falsely grouping LRGs in high-density LRG regions into high $E(B-V)$ bins.
@@
Factor 3 is predicted by Obiwan, since Obiwan simulates the dimming effect from dust extinction. As we discussed in Section 7.1, galaxies with different apparent magnitudes are affected by photometric scattering differently. Faint galaxies are more sensitive towards changes in the depth of images. Here $E(B-V)$ can generate an effect similar to depth by making galaxies fainter. Given its similarity with depth, we cannot see an isolated effect for Factor 3 on DESI LRGs. However, with Obiwan image simulations, we know that these effects are simulated as a whole. If we look at the $E(B-V)$ systematics trend in Figures 12 and 13, we see that the difference in systematics trend of Obiwan LRGs and DESI LRGs are larger in the NGC than the SGC, which is evidenced by $\chi^{2}$ synthetic. This value is 100.5 in NGC and 27.7 in SGC. This indicates that a higher proportion of the $E(B-V)$ trend in the SGC is contributed by phenomena that are simulatable by Obiwan. Factor 3 is one of these phenomena.
@@
We take a deeper look at these simulatable phenomena. We noticed that $E(B-V)$ in the SGC is negatively correlated with psfdepth with a correlation coefficient of about $-0.24$. While in NGC it is positively correlated with a correlation coefficient of about $0.14$. The difference in the two regions is due to the survey design strategy. For the DECaLS data, the DECaLS team adjusted the exposure times to try to reach uniform depth after dust extinction [76]. The reported psfdepth are in observed magnitudes. Thus, the survey goes to deeper psfdepth where the E(B-V) is larger. The positive correlation is expected in the DECaLS area. The SGC has more DES data, and DES did not adopt the depths in the same way as DECaLS. As lower psfdepth leads to higher relative LRG density, this partially explains why we see a stronger positive trend in Obiwan LRGs in the SGC than in the NGC. Due to its close relation with depth, this trend is also related to the magnitude of LRGs. In Fig. 19, we split the sample into faint LRGs and bright LRGs in the same way as in Section 7.1. We see some trend in the SGC faint sample in Obiwan LRGs, while the trend is much weaker in the SGC bright sample. Since faint ones are more influenced by photometric scattering, this behavior is expected. In NGC, given the positive correlation between psfdepth and $E(B-V)$, which cancels out with each other, the trend is much weaker. We can compare the $E(B-V)$ trend of bright DESI LRGs and faint DESI LRGs by looking at the grey and black curve on the left side of Fig. 19. We see that the difference is larger in the SGC than in the NGC. These all indicate that compared to the NGC, the $E(B-V)$ systematics trend in the SGC has a higher portion contributed by Obiwan simulatable phenomena that could be magnitude dependent. As the overall amplitude of the SFD $E(B-V)$ trend is similar in the NGC and the SGC, this suggests that the contributions of factors 1 and 2 (unknown Large-Scale-Structure systematics, including the persistence of CIB contamination) are larger in the NGC than in the SGC. This result is consistent with the stellar-reddening-based map in Fig. 18. The stellar-reddening-map-based $E(B-V)$ trend (green) which is devoid of CIB is more similar to the SFD-map-based $E(B-V)$ trend (blue) from the SFD map in the SGC than in the NGC.
Finally, we want to reiterate the fact that we do not have a “truth map” for $E(B-V)$, which introduces a great challenge to any analysis that involves galaxy samples. Indeed, the two alternative maps we use here show different results, as shown in the green and orange curves in Figure 18. The CSFD map only corrects for the contributions of CIB in the SFD map, so any other Large-Scale-Structure systematics in the SFD map is not recovered by CSFD. Overall, the CSFD $E(B-V)$ map shows a trend more similar to the SFD $E(B-V)$ map than the Mudur+23 map. However, both maps suggest a smaller trend than the SFD $E(B-V)$ map. In the NGC low $E(B-V)$ region in particular, the two alternative maps exhibit a weaker trend compared to the $E(B-V)$ SFD map, suggesting that CIB contamination is likely a main contributing factor to the observed SFD $E(B-V)$ trend in this region.
@@
Figure 19: Systematics trend in $E(B-V)$ when we separate the sample into two subsamples: One bright sample with z band magnitude < 20.5, One faint sample with z band magnitude > 20.5. Blue curves are made with Obiwan-LRGs and black curves are made with real LRGs. The grey curve on the left plots is the same black curve on the right plots, and they are plotted for a better visual comparison. We see some level of trend recovery in SGC faint sample from Obiwan-LRGs, meaning part of the trend is due to simulatable phenomena by Obiwan. The rest does not have a good matching, indicating the trend in real LRGs is likely due to CIB contamination. The result here supports the observation in Figure 18, where we see more $E(B-V)$ systematics trend on CIB-independent map (green) in SGC than in NGC.
7.3 Further Investigation of Trends With PSF Size
Point Spread Function refers to the spread of source profiles due to the Earth’s atmosphere. In the observed images, it makes galaxies more spread out. In Obiwan simulations, we assume the Point spread function (PSF) model is perfect. The pipeline for PSF estimation is PSFEx [77]. This is the same pipeline used in the Dark Energy Survey (DES) Y1 analysis [78] [79]. In DES they saw a bias between the PSF size and the size of stars. They also saw a spatial variation from the fitting residuals. Given the similarity of these two surveys, similar systematics could appear in the fitting of Legacypipe.
Background subtraction is a process of subtracting the background variation in images. The Legacypipe code masks out regions containing sources and fits a spline function to the rest of the images. In regions with larger PSF size, the subtraction could be less accurate because more regions need to be masked from the background estimation, due to all sources having their light more spread in the image. Regions with low depth can also result in more uncertainty in background subtraction. For background subtraction, we currently use the original background determination by Legacypipe as truth. Obiwan injects galaxies on images that are already background subtracted. Possible residuals or over-subtractions in the background subtraction is not replicated in Obiwan.
@@ 7.3 Further Investigation of Trends With PSF Size
The combination of low depth and high psfsize are the regions that are most likely to have false fluctuations that are not simulated by Obiwan. We see that Obiwan LRGs predict that the LRG target density will generally increase with the psfsize, and if anything, the opposite seen in DESI LRGs (though the trends are consistent with the null expectation). To locate the problem, we first split the sample into faint LRGs and Bright LRGs in the same way as in section 7.1. We find that the discrepancy is mainly contributed by faint LRGs. We further split faint LRGs into high depth region with psfdepth_z $>23.5$ and low depth region with psfdepth_z $<23.5$. We find that the trend is mainly contributed from the low-depth region.
@@
Figure 20 shows the trend of DESI LRGs and Obiwan LRGs with these cuts. They are composed of 19.5% of all LRGs in the SGC. We see that the trend of DESI LRGs goes down first. And then it goes up for psfsize_g and psfsize_z. However, Obiwan LRGs does not have a noticeable downward trend, while it does have an upward trend at large psfsize. Plots in Figure 21 indicate the reason behind the mismatch. Here we split the sample to 4 regions by psfsize_z and we then plot their histograms, binning in $z$ magnitude. We normalize these histograms with the random galaxies in the same regions. Using this procedure, we can compare the relative densities of each region, as a function of the $z$ band magnitude. The plot for Obiwan LRGs is produced in the same way as DESI LRGs. The variation we see in DESI LRGs around $z$ band magnitude 21 is not replicated in Obiwan. This discrepancy suggests signs that there is something not simulated by Obiwan, and it is contributed by the PSF. We further study this phenomenon in the following section 7.3.1.
@@
7.3.1 PSF error origination
We first test which LRG color cuts have large scattering among the 4 equations 4.14.4. In Figure 22, We use COSMOS Repeats set 0 as a reference set. For any set $i$ in sets 1 through 9, we compute the number counts of “Loss” and “Gain”. “Loss” is defined as sources passing LRG color cut in reference set 0, but not in set $i$. We consider them “lost LRGs” in set $i$. For each color cut, we compute the number counts of “lost LRGs” that do not pass this cut. “Gain” is the opposite of “Loss”: They are the LRGs that pass the LRG color cut in set $i$, but not in set 0. We consider set $i$ “gained” these LRGs. We check the number counts of these “gained LRGs” that do not pass a certain color cut in set 0. Cuts with high number counts in both the “loss” and “gain” plots are the dominant power to produce photometric scattering, and thus potential density fluctuations. Here we see that cut 2 (equation 4.2) and cut 4 (equation 4.4) have this dominant power in producing density fluctuation. Thus, we focus on these two cuts in the following analysis.
Figure 25: Median $\Delta$fiber flux between the COSMOS Repeats set $i$ ($i$ from 1 to 9) catalogs and the COSMOS Deep catalog, with the COSMOS Deep catalog having a fiber $z$ magnitude range of 21.4 to 21.9. Orange is for all sources while blue is for sources with Sérsic  index >=4. We choose this range because this region is sensitive to LRG color cut 4.2. We see that with an increase of seeing, flux measured in the COSMOS Repeats set is statistically fainter. The trend is prominent for more extended sources. This could contribute to the reason for fewer LRGs observed in regions with worse seeing.
In all, we conclude that we find signs of PSF errors not simulated by Obiwan. This error shifts the statistical distribution of sources to a slightly fainter flux in regions of worse seeing. This shift induces a loss in the number density of LRGs. It could also be the underlying reason for the density dispersion of LRGs around mag z = 21 at different seeing that we see in Figure 21.
7.3.2 Discussion on non-simulated PSF error
Two phenomena are going on in Figure 21. The first phenomenon is that with an increase of psfsize z, there is an increase in the LRG density at the faintest magnitudes, $z>21.3$. This phenomenon is replicated in Obiwan LRGs. This phenomenon is likely because there is more photometric scattering at high psfsize z as these regions typically have lower depth. Sources in this mag z > 21.3 region with the fiber mag z < 21.7 cuts are compact sources. According to results from Obiwan, their true color is more likely to be outside LRG color selection boundaries. The second phenomenon is what we described in section 7.3.1, which causes variation at around mag z = 21. The sources lost due to this phenomenon are likely to have a more extended shape, because they have a larger difference between mag z and fiber mag z.
The sources gained in the first phenomenon and the sources lost in the second phenomenon are not likely to have a large overlap, because of their difference in morphology. Figure 26 shows the COSMOS Deep fiber magnitude and half-light radius of LRGs selected only in COSMOS Repeats set 0 or set 9. Set 0 LRGs are statistically brighter and have a more extended profile than set 9 LRGs. This difference validates our claim that LRGs have a slightly different population at different seeing. For DESI LRGs, we may not see a large density variation in seeing as the “gained” and “lost” in these two phenomena add up. However, it could pose potential risks in cosmological analysis as the two populations of LRGs have different sample statistics. Compact galaxies have less star-forming processing going on than extended galaxies. Even if they are similar in photometric color, they may differ in stellar mass, which results in a difference in galaxy bias.
There may not be a straightforward way to quantify this issue. For DESI LRGs, we will eventually obtain the spectrum of all LRGs, and knowing the redshift of each source will help better address this issue. However, imaging surveys like LSST or DES may encounter similar problems. Our study motivates the development of a synthetic source injection pipeline that includes the effects of background subtraction errors and PSF model errors. With background subtraction and PSF errors properly simulated, we would be able to replicate the behavior in real data. This could give us a better understanding of galaxy density fluctuation so that we would be able to correct for these trends accordingly.

### v2

@@ 7.2 Galactic Extinction
Various maps have been developed to correct for galactic extinction. As detailed in [69], these maps can be divided into three categories: Thermal emission-based maps, e.g., [66], stellar reddening-based maps, e.g., [70, 71], and HI-based maps, e.g., [72]. The Galactic extinction is typically mapped as $E(B-V)$, in magnitudes, and a correction to apply to any measured magnitude is a constant coefficient [73] for the particular wavelength band, multiplied by the $E(B-V)$ at the given celestial coordinate. The SFD ([66]) map is the most commonly used map for the application of extinction corrections in current galaxy surveys, including the DESI LRGs targets. It is thermal-based, and the data are directly collected from dust emission. This gives it high accuracy. However, it suffers from extragalactic imprints. When stacking the SFD $E(B-V)$ value around reference objects, [69] see that the average $E(B-V)$ value correlates with the redshift of these objects, indicating leaks of Cosmic Infrared Background (CIB) signal. The stellar-reddening-based maps measure dust emission on a different wavelength range, in which CIB contamination is negligible. These maps measure $E(B-V)$ by comparing stellar spectra to templates while fitting for the $E(B-V)$ value at the same location. The accuracy of this method is limited by the number of usable stars at any given location. Moreover, the zero-point calibration uncertainties from star catalogs used for stellar-reddening-based $E(B-V)$ maps could bias the measurements for these maps [74]. However, they are free from CIB leaks. HI-based maps trace the hydrogen column density, correlated with $E(B-V)$. HI-based maps are not strictly $E(B-V)$ maps. Due to the scope of this work, we did not use HI-based maps in our test.
Observations of distant galaxies are affected by $E(B-V)$. If we have a “truth map” for $E(B-V)$ (and per-band coefficients), it would not be a big problem, as the needed corrections to the photometry would be known. However, we do not have such a “truth map”. When we detect a correlation of our galaxies’ relative density with the $E(B-V)$ map, we are not certain how the error in our $E(B-V)$ map plays a role in such correlation.
We will consider three potential factors that could contribute to the trends we observe between $E(B-V)$ and DESI LRGs:
@@
    The $E(B-V)$ values are correlated with the true background galaxy density due to some systematics in the $E(B-V)$ map, e.g., the CIB leaks. For LRGs, their own dust can absorb light and re-emit it in the infrared, contributing to the CIB. Thus, regions with a higher density of LRGs will have higher CIB and a positive trend should be expected between the LRG density and an $E(B-V)$ map with CIB contamination.
@@
Factors 1 and 2 are not recovered by Obiwan because Obiwan assumes no error in the $E(B-V)$ map, while factor 3 should be automatically included during Obiwan’s image simulation procedure. In DESI LRGs, we observe a strong correlation of the relative number density of LRGs with the $E(B-V)$ SFD map, as shown in Figures 7 and 8. Meanwhile, Obiwan LRGs exhibit a weaker correlation. Under our assumption, the weaker correlation in Obiwan LRGs is due to not simulating factors 1 and  2 in the Obiwan simulation.
To test the impact of the three factors on the LRG density vs. $E(B-V)$ trend, we redo the DESI LRGs target selection based on two alternative maps. One of the alternative maps is the stellar-reddening-based map produced by [70], which we label ‘Mudur+23’. This map derives $E(B-V)$ map based on data from Pan-STARRS1 [75] and 2MASS [76] along with Gaia parallaxes [77]. The other alternative map is the Corrected-SFD (CSFD) map produced by [78]. It cross-correlated the SFD map with the galaxy clustering signal and separated the SFD map into two maps: the CIB residual map and the CSFD map. The CSFD dust map has a much lower residual CIB contamination than SFD. However, the correction is CIB only. Non-CIB systematics are not included in the CSFD map. Compared to the stellar reddening-based map, the $E(B-V)$ values on the CSFD map share more similarity with the SFD map. We define LRGs selected with these maps real LRGs, since they are all real galaxies we observe. When using an alternative $E(B-V)$ map, real LRGs is not always identical to DESI LRGs. When we use the alternative $E(B-V)$ maps, the (extinction-corrected) magnitudes of the galaxies change. When we apply LRG target selection to these galaxies, we obtain a sample slightly different from DESI LRGs. Meanwhile, if we use SFD map in LRG target selection, real LRGs are identical to DESI LRGs.
@@
Figure 12: Density fluctuation of DESI LRGs based on different maps in the NGC (left) and SGC (right) region. The target selection process is also altered based on different $E(B-V)$ maps. The blue line is based on the SFD map [66], the green line (labeled ‘Mudur+23’) is based on the stellar reddening map from [70], the orange line is based on the Corrected-SFD (CSFD) map [78], the red line uses LRGs selected by the SFD map, but the binning is computed with the same stellar reddening map as the green line. For LRGs selected and plotted against SFD maps, we see a clear trend. In the stellar reddening map, NGC and SGC are slightly different. In NGC, we do not see any trend. In SGC, we see a trend that is weaker than the SFD map. The difference is also seen in Obiwan LRGs seen in Figure 13. In the CSFD map, this trend is similar to the SFD map. In the lower $E(B-V)$ region, the trend is weaker. The fact that the red lines exhibit a larger trend than the green line suggests that some of the $E(B-V)$ systematics trend is introduced by target selection, which uses the $E(B-V)$ to determine the colors of galaxies.
Factor 1 can be seen by comparing the red and green curves. The two samples are binned with the same stellar-reddening-based map. The red sample is DESI LRGs that uses the SFD map to determine their magnitudes and perform target selection. The green curve uses a stellar-reddening-based map for a similar procedure. Thus, the difference between these two curves is a result of the change in target selection. We find that 2.0% of LRGs are selected with the stellar-reddening based map, and not selected by the SFD map. 2.6% of LRGs are selected with the SFD map, but not selected by the stellar-reddening based map.
These two curves in Figure 12 show how LRG sample difference results in different systematics trends. The red curve has a larger trend than the green curve, which indicates that the CIB leaks affect the target selection of DESI LRGs.
Factor 2 can be seen by comparing the blue and red curves. The two trends are computed with identical DESI LRGs sample. The only difference lies in the binning. The blue curve is SFD-map-selected LRGs binned with the SFD map. It has a stronger trend. The red curve is the same SFD-map-selected LRGs but binned with the stellar-reddening-based map. It has a weaker trend. The reduction of trend in the red curve compared to the blue curve is partly a result of less CIB contamination in the stellar-reddening-based map. However, it is worth noting that this trend reduction could also result from uncertainties in the Mudur+23 map: If we add some noise to the “$E(B-V)$ truth map”, then some $E(B-V)$ values will be falsely assigned to other bins, resulting in a suppression of the actual trend. It could also result from SFD $E(B-V)$ map uncertainty: as the galaxies are selected with the SFD map, binning with the same map includes relevant unknown systematics.
@@
Factor 3 is predicted by Obiwan, since Obiwan simulates the dimming effect from dust extinction. As we discussed in Section 7.1, galaxies with different apparent magnitudes are affected by photometric scattering differently. Faint galaxies are more sensitive towards changes in the depth of images. Here $E(B-V)$ can generate an effect similar to depth by making galaxies fainter. Given its similarity with depth, we cannot see an isolated effect for Factor 3 on DESI LRGs. However, with Obiwan image simulations, we know that these effects are simulated as a whole. If we look at the $E(B-V)$ systematics trend in Figures 7 and 8, we see that the difference in systematics trend of Obiwan LRGs and DESI LRGs are larger in the NGC than the SGC, which is evidenced by $\chi^{2}$ synthetic. This value is 100.5 in NGC and 27.7 in SGC. This indicates that a higher proportion of the $E(B-V)$ trend in the SGC is contributed by phenomena simulatable by Obiwan. Factor 3 is one of these phenomena.
@@
We take a deeper look at these simulatable phenomena. We noticed that $E(B-V)$ in the SGC is negatively correlated with psfdepth with a correlation coefficient of about $-0.24$. While in NGC it is positively correlated with a correlation coefficient of about $0.14$. The difference in the two regions is due to the survey design strategy. For the DECaLS data, the DECaLS team adjusted the exposure times to try to reach uniform depth after dust extinction [80]. The reported psfdepth are in observed magnitudes. Thus, the survey goes to deeper psfdepth where the E(B-V) is larger. The positive correlation is expected in the DECaLS area. The SGC has more DES data, and DES did not adopt the depths in the same way as DECaLS. As lower psfdepth leads to higher relative LRG density, this partially explains why we see a stronger positive trend in Obiwan LRGs in the SGC than in the NGC. Due to its close relation with depth, this trend is also related to the magnitude of LRGs. In Figure 13, we split the sample into faint LRGs and bright LRGs in the same way as in Section 7.1. We see some trend in the SGC faint sample in Obiwan LRGs, while the trend is much weaker in the SGC bright sample. Since faint ones are more influenced by photometric scattering, this behavior is expected. In NGC, given the positive correlation between psfdepth and $E(B-V)$, which cancels out with each other, the trend is much weaker. We can compare the $E(B-V)$ trend of bright DESI LRGs and faint DESI LRGs by looking at the grey and black curve on the left side of Figure 13. We see that the difference is larger in the SGC than in the NGC. These all indicate that compared to the NGC, the $E(B-V)$ systematics trend in the SGC has a higher portion contributed by Obiwan simulatable phenomena that could be magnitude dependent. As the overall amplitude of the SFD $E(B-V)$ trend is similar in the NGC and the SGC, this suggests that the contributions of factors 1 and 2 (unknown $E(B-V)$ systematics, including the persistence of CIB contamination) are larger in the NGC than in the SGC. This result is consistent with the stellar-reddening-based map in Figure 12. The stellar-reddening-map-based $E(B-V)$ trend (green) which is devoid of CIB is more similar to the SFD-map-based $E(B-V)$ trend (blue) from the SFD map in the SGC than in the NGC. Our findings here are consistent with [81] (particularly Figure 15 in their paper), and we further discuss it in the next paragraph.
In Figure 12, the two alternative maps we use here have different trends, as shown in the green and orange curves. The CSFD map only corrects for the contributions of CIB in the SFD map, so any other systematics in the SFD map is not recovered by CSFD. As discussed in [70], there is a negative trend for star-reddening-based $E(B-V)$ $E(B-V)$ SFD as a function of $E(B-V)$ SFD. This trend supports the conclusion in [81]. In their work, they computed the reddening coefficient with the blue tip of the stellar locus and the measurement is based on the $E(B-V)$ SFD map. They found a dust map normalization difference of 15% between NGC and SGC that may be due to uncertainties in dust temperature correction. The dust temperature map has high uncertainty, particularly in the low $E(B-V)$ region. The angular resolution of the dust temperature map is much lower than the $E(B-V)$ map. As the $E(B-V)$ scales linearly with the dust temperature map, uncertainties in the dust temperature map bias the overall amplitude of $E(B-V)$ rather than the relative fluctuations. This effect generates a trend as a function of the absolute $E(B-V)$ value.
@@
Figure 13: Systematics trend in $E(B-V)$ when we separate the sample into two subsamples: One bright sample with z band magnitude < 20.5, One faint sample with z band magnitude > 20.5. Blue curves are made with Obiwan-LRGs, and black curves are made with real LRGs. The grey curve on the left plots is the same black curve on the right plots, and they are plotted for a better visual comparison. We see some trend recovery in the SGC faint sample from Obiwan-LRGs, meaning part of the trend is due to simulatable phenomena by Obiwan (e.g., $E(B-V)$ in the SGC produces a shallower depth that is not corrected in the observing strategy design). The rest do not match well, indicating that the trend in real LRGs is likely due to the SFD $E(B-V)$ map systematics. The result here supports the observation in Figure 12, where we see more discrepancies between the SFD $E(B-V)$ and alternative maps in NGC than in SGC.
7.3 Investigation of Trends With psfsize
@@
Obiwan LRGs predict that the LRG target density generally increases with psfsize, and if anything, the opposite seen in DESI LRGs (though the trends are consistent with the null expectation). To locate the problem, we first split the sample into faint LRGs and Bright LRGs in the same way as in Section 7.1. The discrepancy is mainly contributed by faint LRGs. We further split faint LRGs into a high-depth region with psfdepth_z $>23.5$ and a low-depth region with psfdepth_z $<23.5$. We find that the trend is mainly contributed from the low-depth region.
@@
Figure 14 shows the trend of DESI LRGs and Obiwan LRGs with these cuts (magnitude $z$ > 20.5, psfdepth_z < 23.5). They are composed of 19.5% of all LRGs in the SGC. The trend of DESI LRGs goes down first. And then it goes up for psfsize_g and psfsize_z. However, Obiwan LRGs does not have a noticeable downward trend, while it does have an upward trend at large psfsize. Plots in Figure 15 indicate the reason behind the mismatch. Here, we split the sample into four regions by psfsize_z, then plot their histograms, binning in $z$ magnitude. We normalize these histograms with the random galaxies in the same regions. Using this procedure, we can compare the relative densities of each region as a function of the $z$ band magnitude. The plot for Obiwan LRGs is produced in the same way as DESI LRGs. The variation we see in DESI LRGs around $z$ band magnitude 21 is not replicated in Obiwan.
The COSMOS repeat sets are ideal sets for pinning down this issue. We perform several tests documented in Appendix E. We eventually concluded that this effect is caused by a correlation of flux bias with psfsize, shown in Figure 25. The trend for this bias is larger for galaxies with Sérsic  index >=>=> =4 (we define them as disk LRGs). Because of this dependency with galaxy morphology, disk LRGs have a lower flux at high psfsize_z, so they are less likely to be selected as LRGs due to the color cut 4.2. Consequently, we see less disk LRGs in regions with large psfsize. This effect is further propagated to DESI LRGs after spectroscopic redshift cuts. As disk LRGs have a star-forming disk, they have stronger OII emission lines. We use the DESI LRGs with spectroscopic redshift from the iron catalog. We compute the median OII flux for each psfsize_z bin. This gives us an estimate of the median intensity of OII flux for each psfsize_z bin. In Figure 16, we found noticeable trends in all three redshift bins. The median OII intensity decreases with an increase in psfsize_z. We do not see similar trends in the north (BASS/MzLS region). As we have not thoroughly tested the north in this work, we do not know why similar trends are not seen in the north.
@@
Figure 16: The median flux of OII emission line for LRGs at three redshift bins in different psfsize_z range, divided by the overall median OII flux. Blue is for the redshift range 0.4-0.6, orange is for the redshift range 0.6-0.8, and green is for the redshift range 0.8-1.1. The trend here originates from a higher loss in flux for disk LRGs at high psfsize_z
7.3.1 Possible root causes
Knowing where the discrepancy originates will help us develop treatments for this issue, so we did some further investigation. Our main suspects of root causes are background subtraction and PSF uncertainty.
We currently use the original background determination by Legacypipe as truth. Obiwan injects galaxies on images that are already background subtracted. Possible residuals or over-subtractions in the background subtraction are not replicated in Obiwan. The background subtraction procedure in the Legacypipe code masks out regions containing sources and fits a spline model [82] to the rest of the images.
In the case of PSF uncertainty, it is a widely studied subject. In Dark Energy Survey (DES) analysis [83] [84], people saw a bias between the PSF size and the size of stars (Brighter-Fatter effect). They also saw a color dependency in the fitting residuals (Differential Chromatic Refraction). The pipeline (PSFEx[85]) used for PSF determination in Legacysurvey is similar to DES (PIFF[86]), so similar systematics could also be present in Legacysurvey.
We design a test to separate the impacts of background subtraction uncertainty and PSF uncertainty. We apply a Gaussian filter to each individual image (CCD) to make them reach the same psfsize. Then, we measure the 7" aperture flux of the same LRGs on different CCDs. A 7" aperture is large enough to contain most of the flux from LRGs. The 7" aperture flux is independent of any fitting methods or determination of PSF. On each CCD, we record the measured flux and the psfsize value before Gaussian filtering. These data are later used to test the correlation between the measured flux and the psfsize. We perform the same test with synthetic Obiwan LRGs on real images and also on fully synthetic images with Gaussian background noise for comparison.
We select all pairs of images with a difference in psfsize_z > 0.12". With these data pairs, we found that the 7" aperture flux for the images with a larger PSF is 0.19 nanomaggies lower than the images with a smaller PSF. The equivalent value measured from the synthetic Obiwan LRGs with synthetic images is less than $\sim$10-3 nanomaggies. For Obiwan LRGs injected on real images, this number is 0.09 nanomaggies. Overall, we conclude that the background on a larger psfsize is lower than that on images with smaller psfsize. The flux loss in DESI LRGs is $\sim$0.1 nanomaggies larger than the synthetic Obiwan LRGs. However, we do not find the flux loss for Disk LRGs different from Compact LRGs. The conclusion here applies to 3 optical ($g$, $r$ $z$) bands.
Our test on 7" aperture flux suggests that there could be some flux bias induced by background subtraction. In regions with larger psfsize, it becomes more difficult to distinguish between signals from real galaxies and signals from the background, as the galaxies on images are more spread out. The signal-to-noise is lower on each pixel. The procedure of source fitting uses a set of effective pixels. The fitting of Disk LRGs requires more pixels, so it is more sensitive to the cumulative effect of the flux bias from the background subtraction.
On the other hand, uncertainties from PSF determination may not be a dominant factor here, as our PSF-insensitive test is already sufficient to explain the trends with psfsize.